The square root of the sum of the square of a number increased by 5 added to the number is equal to 5. Find the number
Answer by fractalier(6550)

Writing the equation carefully, we get
sqrt(x^2 + 5) + x = 5
Then bring the x to the other side and square both sides
sqrt(x^2 + 5) = 5 - x
x^2 + 5 = 25 - 10x + x^2
Subtract x^2...
5 = 25 - 10x
-20 = -10x
x = 2
Normally you check any kind of root problem...this solution works...